OpenStudy (anonymous):

Given the linear equation y=-7/8x+2, find the y-coordinates of the points(-8),(0), and (8). Plot those points and graph the linear equation.

OpenStudy (anonymous):

the points they gave you were for x so plug in the equation and you will get the value of y i will do one for you - x = -8 y = -7/8(x) +2 y = -7/8 (-8) + 2 y = 7 + 2 y = 9